# DeepLag: Discovering Deep Lagrangian Dynamics for Intuitive Fluid Prediction

### v3

@@ 4.4 Model Analysis
To verify the effectiveness of detailed designs in DeepLag, we conduct exhaustive ablations in Table 5. In our original design, we track 512 particles (around 3% of Eulerian grids).
Table 5: Ablations on Bounded Navier-Stokes, including reducing tracking particles of the first layer ($M_{1}$), removing Lagrangian-guided feature evolving $\operatorname{LagToEu}(\cdot)$, Eulerian-conditioned particle tracking $\operatorname{EuToLag}(\cdot)$ and learnable particle sampling strategy.
Design Relative L2 ($\downarrow$) Decrease
DeepLag ($M_{1}=512$) 0.0543 -
$M_{1}=128$ 0.0559 2.9%
$M_{1}=256$ 0.0553 1.8%
$M_{1}=768$ 0.0547 0.7%
w/o $\operatorname{LagToEu}(\cdot)$ 0.0556 2.4%
w/o $\operatorname{EuToLag}(\cdot)$ 0.0547 0.7%
w/o Learnable Sampling 0.0552 1.6%
The experiments indicate that further increasing the number of particles yields marginal performance improvements. Therefore, we opt for $M_{1}=512$ to strike a balance between efficiency and performance. In addition, we can conclude that all components proposed in this paper are indispensable. Especially, the lack of interaction between the Eulerian and Lagrangian space will cause a severe drop in accuracy, highlighting the dual cross-attention that exploits Lagrangian dynamics has a positive influence on the evolution of Eulerian features. Besides, rather than uniformly sampling particles, sampling from a learnable probability matrix also provides an upgrade (refer to Appendix G for visual results). The above results provide solid support to our motivation in tracking essential particles and utilizing Eulerian-Lagrangian recurrence, further confirming the merits of our model.
@@ 5 Conclusions and Limitations
To tackle intricate fluid dynamics, this paper presents DeepLag by introducing the Lagrangian dynamics into Eulerian fluid, which can provide clear and neat dynamics information for prediction. A EuLag Block is presented in the Eulerian-Lagrangian Recurrent Network to utilize the complementary advantages of Eulerian and Lagrangian perspectives, which brings better particle tracking and Eulerian fluid prediction. DeepLag excels in complex fluid prediction with an average improvement of nearly 10% across three carefully selected complex benchmarks, even in 3D fluid, and can also provide interpretable evidence by plotting learned Lagrangian trajectories. However, the number of tracking particles in DeepLag is appointed as a fixed hyperparameter that needs to be adjusted for specific scenarios, leaving space for future exploration.
@@ A.3 EuLag Block
$\begin{split}\displaystyle\hat{\mathbf{u}}_{t+1}^{l}&={\mathbf{u}}_{t}^{l}+\operatorname{LagToEu}({\mathbf{u}}_{t}^{l},{\mathbf{h}}_{t}^{\prime l},{\mathbf{h}}_{t}^{\prime l})\\ &={\mathbf{u}}_{t}^{l}+\operatorname{FFN}(\operatorname{LagToEu\_Attn}({\mathbf{u}}_{t}^{l},{\mathbf{h}}_{t}^{\prime l},{\mathbf{h}}_{t}^{\prime l}))+\operatorname{LagToEu\_Attn}({\mathbf{u}}_{t}^{l},{\mathbf{h}}_{t}^{\prime l},{\mathbf{h}}_{t}^{\prime l}),\ \text{$l$ from $1$ to $L$},\end{split}$ (15)
@@ A.4 Metrics
Appendix B Comparison Between DeepLag and Neural ODE-based Methods
DeepLag differs significantly from ODE-based approaches like Neural ODE [4], as highlighted in the comparison Table 7:
Table 7: Comparison between DeepLag and ODE-based models
@@ C.1 Bounded Navier-Stokes
Our data generation method involves running simulations for over $10^{5}$ steps after setting the initial conditions of the flow field. We then randomly select a starting time step and extract several frames as an example, which are further randomly divided into training and testing sets. Due to the highly unsteady nature of the flow field, the flow patterns observed by the model appear significantly different.
@@ C.2 Ocean Current
The procedures to make this dataset are as follows. First, we downloaded daily sea reanalysis data [5] from 2011 to 2020 provided by the ECMWF and selected five basic variables on the sea surface to construct the dataset, including velocity, salinity, potential temperature, and height above the geoid, which are necessary to identify the ocean state. Then, we crop a $180\times 300$ sub-area on the North Pacific from the global record, corresponding to a 375km×$\times$625km region. In total, this dataset consists of 3,653 frames, where the first 3000 frames are used for training and the last 600 frames are used for testing. The training task is to predict the future current of 10 days based on the past 10 days’ observation, after which we performed 30 days of inference with the trained model to examine the long-term stability of DeepLag.
@@ Appendix D Training Curves
We provide training curves on Bounded Navier-Stokes, Ocean Current and 3D Smoke datasets in Figure 8. We can observe that DeepLag presents favorable training robustness and converges the fastest on the Bounded Navier-Stokes dataset.
@@ Appendix F ACC Metric on Ocean Current
In meteorology, directly calculating the correlation between predictions and ground truth may obtain misleadingly high values because of the seasonal variations. To subtract the climate average from both the forecast and the ground truth, we utilize the Anomaly Correlation Coefficient to verify the forecast and observations. Moreover, since the observation grids are equally spaced in latitude, the size of the different grids is related to the latitude and thus, we calculate the latitude-weighted Anomaly Correlation Coefficient, which can be formalized as:
@@
Notably, DeepLag also excels in the ACC metric as shown in Table 11, which can better quantify the model prediction skill. As shown in the timewise ACC curve, DeepLag consistently achieves the highest ACC and holds more significant advantages in long-term prediction. Since ACC is calculated with long-term climate statistics, further improvements become increasingly challenging as it increases, which highlights the value of DeepLag.
@@ Appendix G Visual Results for learnable sampling
Figure 9: Visualization of the pointwise variance of vorticity (middle) and the sampling probability distribution (right) learned by DeepLag. For a better view, we plot the $\log(\mathbf{S})$ here.
To demonstrate the effectiveness of our learnable probability, we visualize its distribution $\mathbf{S}$ with respect to the pointwise variance of vorticity, which is directly proportional to the local complexity of fluid dynamics, in Figure 9. It is evident that our sampling module tends to prioritize regions with higher dynamical complexity, such as the wake flow and Karman vortex [50] behind the pillar. The showcase in Figure 1 also demonstrates that the tracked particle can well present the dynamics of a certain area. This observation underscores that our design effectively guides the model to track the most crucial particles, enhancing its performance.
@@ I.3 Examination on the turbulent kinetic energy spectrum
In the field of fluid mechanics, simulation results that better adhere to intrinsic physical laws are sometimes more valuable than those with smaller pointwise errors, often reflected in frequency domain analysis. To validate this point and to measure long-term forecasting ability, we introduced a metric on time-averaged turbulent statistics, namely the turbulent kinetic energy spectrum (TKES). Specifically, we computed the MAE and RMSE of TKES for the Bounded Navier-Stokes dataset, which exhibits the most prominent turbulent characteristics, as shown in Table 14, and plotted the line graphs of the error of wave number and TKE, as depicted in Figure 26. It can be observed both numerically and visually that DeepLag consistently outperforms various baselines to different extents.
Table 14: MAE and RMSE of TKES on Bounded Navier-Stokes.

### v4

@@ 4.4 Model Analysis
To verify the effectiveness of detailed designs in DeepLag, we conduct exhaustive ablations in Table 5. In our original design, we track 512 particles (around 3% of Eulerian grids) in total on 4 hierarchical spatial scales with a latent dimension 64 in the Lagrangian space.
Table 5: Ablations of (a) module removing and (b) hyperparameter sensitivity on Bounded Navier-Stokes, where (a) includes w/o $\operatorname{LagToEu}$: removing Lagrangian-guided feature evolving, w/o $\operatorname{EuToLag}$: removing Eulerian-conditioned particle tracking and w/o Learnable Sampling: removing learnable particle sampling strategy, and (b) includes #Particle ($M_{1}$): adjusting the number of tracking particles of the first layer, #Scale ($L$): adjusting the number of spatial scales and #Latent ($C_{1}$): adjusting the number of latent dimension of the first layer. We use the control variable method, when one hyperparameter is changed, the values of the others are all original (ori). Ablations of (c) attention swapping on Bounded Navier-Stokes (2D) and 3D Smoke (3D), where the order of $\operatorname{EuToLag}$ and $\operatorname{LagToEu}$ blocks is swapped. Original and Swapped Relative L2 are reported.
(a) Module Removing
Design Relative L2 ($\downarrow$)
DeepLag 0.0543
w/o $\operatorname{LagToEu}(\cdot)$ 0.0556
w/o $\operatorname{EuToLag}(\cdot)$ 0.0547
w/o Learnable Sampling 0.0552
(b) Hyperparameter Sensitivity
#Particle Relative L2 ($\downarrow$) #Scale Relative L2 ($\downarrow$) #Latent Relative L2 ($\downarrow$)
$128$ 0.0559 $1$ 0.0789 $16$ 0.0656
$256$ 0.0553 $2$ 0.0658 $32$ 0.0594
$512$(ori) 0.0543 $4$(ori) 0.0543 $64$(ori) 0.0543
$768$ 0.0547 $5$ 0.0554 $128$ 0.0614
(c) Attention Swapping
Data Original ($\downarrow$) Swapped ($\downarrow$)
2D 0.0543 0.0545
3D 0.0378 0.0378
@@
The experiments indicate that further increasing the number of particles, scales, or latent dimensions yields marginal performance improvements. Therefore, we opt for these values to strike a balance between efficiency and performance. In addition, we can conclude that all components proposed in this paper are indispensable. Especially, the lack of interaction between the Eulerian and Lagrangian space will cause a severe drop in accuracy, highlighting the dual cross-attention that exploits Lagrangian dynamics has a positive influence on the evolution of Eulerian features. Besides, rather than uniformly sampling particles, sampling from a learnable probability matrix also provides an upgrade (refer to Appendix G for visual results). When swapping the positions of the $\operatorname{EuToLag}$ and $\operatorname{LagToEu}$ blocks, the minimal performance change in both 2D and 3D benchmarks suggests the equivalence of the two perspectives, and the flow of information between them is bidirectional and insensitive to the order, underscoring the robustness of our approach. The above results provide solid support to our motivation in tracking essential particles and utilizing Eulerian-Lagrangian recurrence, further confirming the merits of our model.
@@
Table 6: Experiments of (a) model efficiency alignment on 3D Smoke and (b) high-resolution data on Bounded Navier-Stokes. To compare with an aligned efficiency, we add more convolutional layers into the standard U-Net and increase the latent dimension. In the high-resolution simulation, we trained a new DeepLag model on a newly generated 256$\times$256 Bounded Navier-Stokes dataset. Model parameters (#Param), GPU memory (Mem) and running time per epoch (Time) are reported.
(a) Model Efficiency Alignment
Model #Param Mem Time Relative L2 ($\downarrow$)
U-Net-scale 336M 13672MB 601s/ep NaN
DeepLag 19M 12112MB 845s/ep 0.0378
(b) High-resolution Data
Resolution Mem Time Relative L2 ($\downarrow$)
$128\times 128$ 5420MB  1150s/ep 0.0543
$256\times 256$ 13916MB  1300s/ep 0.0514
Comparison under aligned efficiency
As aforementioned, U-Net also presents competitive performance and efficiency in some cases. To carry out a fair comparison between baselines and DeepLag, we examine them on 3D Smoke, which scales up U-Net to have a running time similar to DeepLag as in Table 6(a). The results show that too many parameters overwhelm small baselines at the lower left corner in Figure 7 like U-Net, indicating that they have a shortcoming in scalability.
Generalization analysis
To verify the generalizing ability of DeepLag on larger and new domains, we ran tests on high-resolution (HR) data as in Table 6(b) and unseen boundary conditions (BC) as in Appendix J, respectively. The HR simulation on the $256\times 256$ Bounded Navier-Stokes shows that the finer the data are, the more accurate our DeepLag is, which still has a 17% promotion w.r.t U-Net on relative L2 (DeepLag: 0.051, U-Net: 0.060). The comparison of GPU memory and running time per epoch also confirms that the increase of the space complexity is sublinear and the time complexity is minor, underscoring the scalability of DeepLag. For BCs, we ran a zero-shot test with the old model checkpoint on a newly generated Bounded N-S dataset, which has obstacles of different numbers, positions, and sizes. DeepLag still has a 7% promotion w.r.t the best baseline, U-Net, on relative L2 (DeepLag: 0.203, U-Net: 0.217). The visual comparison between the two models further shows that DeepLag adaptively generalizes well on unknown and more complex domains.
@@ 5 Conclusions and Limitations
To tackle intricate fluid dynamics, this paper presents DeepLag by introducing the Lagrangian dynamics into Eulerian fluid, which can provide clear and neat dynamics information for prediction. A EuLag Block is presented in the Eulerian-Lagrangian Recurrent Network to utilize the complementary advantages of Eulerian and Lagrangian perspectives, which brings better particle tracking and Eulerian fluid prediction. DeepLag excels in complex fluid prediction with an average improvement of nearly 10% across three carefully selected complex benchmarks, even in 3D fluid, and can also provide interpretable evidence by plotting learned Lagrangian trajectories. However, the number of tracking particles in DeepLag is appointed as a fixed hyperparameter that needs to be adjusted for specific scenarios, and missing Lagrangian supervision in the data leads to a lack of judgment on learned Lagrangian dynamics, which leave space for future exploration. That said, results show that the performance always tends to incline as we simply scaling up, therefore the few hyperparameters for tuning depend more on the limitation of computing resources rather than blind searching.
Acknowledgments and Disclosure of Funding
This work was supported by the National Natural Science Foundation of China (U2342217 and 62022050), the BNRist Project, and the National Engineering Research Center for Big Data Software.
@@ A.2.2 Multiscale architecture
For position embedding, we concatenate two additional channels to the input (three for 3D Smoke), representing normalized $(x,y)$ coordinates (or $(x,y,z)$ for 3D Smoke).
@@ A.3 EuLag Block
$\begin{split}\displaystyle\tilde{\mathbf{u}}_{t+1}^{l}&={\mathbf{u}}_{t}^{l}+\operatorname{LagToEu}({\mathbf{u}}_{t}^{l},{\mathbf{h}}_{t}^{\prime l},{\mathbf{h}}_{t}^{\prime l})\\ &={\mathbf{u}}_{t}^{l}+\operatorname{FFN}(\operatorname{LagToEu\_Attn}({\mathbf{u}}_{t}^{l},{\mathbf{h}}_{t}^{\prime l},{\mathbf{h}}_{t}^{\prime l}))+\operatorname{LagToEu\_Attn}({\mathbf{u}}_{t}^{l},{\mathbf{h}}_{t}^{\prime l},{\mathbf{h}}_{t}^{\prime l}),\ \text{$l$ from $1$ to $L$},\end{split}$ (17)
@@ A.4 Metrics
Appendix B Comparison Between DeepLag and Classical ML Methods
DeepLag differs significantly from classical ML approaches like Neural ODE [4], as highlighted in the comparison Table 8:
Table 8: Comparison between DeepLag and classical ML model
@@ C.1 Bounded Navier-Stokes
Our data generation method involves running simulations for over $10^{5}$ steps after setting the initial conditions of the flow field. We then randomly select a starting time step and extract several frames as an example, which are further randomly divided into training and testing sets. The positions of the cylinders are fixed, but the initial condition varies in different samples, which can simulate a scenario like bridge pillars in a torrential river. Due to the highly unsteady nature of the flow field, the flow patterns observed by the model appear significantly different.
@@ C.2 Ocean Current
The procedures to make this dataset are as follows. First, we downloaded daily sea reanalysis data [5] from 2011 to 2020 provided by the ECMWF and selected five basic variables on the sea surface to construct the dataset, including velocity, salinity, potential temperature, and height above the geoid, which are necessary to identify the ocean state. Then, we crop a $180\times 300$ sub-area on the North Pacific from the global record, corresponding to a 375km×$\times$625km region. In total, this dataset consists of 3,653 frames, where the first 3000 frames are used for training, and the last 600 frames are used for testing. The training task is to predict the future current of 10 days based on the past 10 days’ observation, after which we performed 30 days of inference with the trained model to examine the long-term stability of DeepLag.
@@ Appendix D Training Curves
We provide training curves on Bounded Navier-Stokes, Ocean Current, and 3D Smoke datasets in Figure 8. We can observe that DeepLag presents favorable training robustness and converges the fastest on the Bounded Navier-Stokes dataset.
@@ Appendix F ACC Metric on Ocean Current
In meteorology, directly calculating the correlation between predictions and ground truth may obtain misleadingly high values because of the seasonal variations. To subtract the climate average from both the forecast and the ground truth, we utilize the Anomaly Correlation Coefficient to verify the forecast and observations. Moreover, since the observation grids are equally spaced in longitude, and the size of the different grids is related to the latitude, we calculate the latitude-weighted Anomaly Correlation Coefficient, which can be formalized as:
@@
Notably, DeepLag also excels in the ACC metric, as shown in Table 12, which can better quantify the model prediction skill. As shown in the timewise ACC curve, DeepLag consistently achieves the highest ACC and holds more significant advantages in long-term prediction. Since ACC is calculated with long-term climate statistics, further improvements become increasingly challenging as it increases, which highlights the value of DeepLag.
@@ Appendix G Visual Results for Learnable Sampling
Figure 9: Visualization of the pointwise variance of vorticity (middle) and the sampling probability distribution (right) learned by DeepLag. We plot the $\log(\mathbf{S})$ here for a better view.
To demonstrate the effectiveness of our learnable probability, we visualize its distribution $\mathbf{S}$ with respect to the pointwise variance of vorticity, which is directly proportional to the local complexity of fluid dynamics, in Figure 9. It is evident that our sampling module tends to prioritize focusing on and sampling particles within regions having higher dynamical complexity, such as the wake flow and Karman vortex [50] near the domain borders and behind the pillar. The showcase in Figure 1 also demonstrates that the tracked particle can well present the dynamics of a certain area. This observation underscores that our design is very flexible at adapting to various complex boundary conditions and can effectively guide the model to track the most crucial particles, enhancing its performance in capturing the fine details in the wake zone where turbulence and vortex form.
@@ I.3 Examination on the turbulent kinetic energy spectrum
In the field of fluid mechanics, simulation results that better adhere to intrinsic physical laws are sometimes more valuable than those with smaller pointwise errors, often reflected in frequency domain analysis. To validate this point and to measure long-term forecasting ability, we introduced a metric on time-averaged turbulent statistics, namely the turbulent kinetic energy spectrum (TKES). Specifically, we computed the MAE and RMSE of TKES for the Bounded Navier-Stokes dataset, which exhibits the most prominent turbulent characteristics, and plotted the line graphs of the error of wave number and TKE, as shown in Table 15. It can be observed both numerically and visually that DeepLag consistently outperforms various baselines to different extents.
Table 15: Turbulent kinetic energy spectrum (TKES) results on the Bounded Navier-Stokes. Both the plot of error of TKES (left) and MAE and RMSE of TKES (right) are presented. A lower TKES error value indicates better performance. Relative promotion is also calculated.